# A Study of Undefined Behavior Across Foreign Function Boundaries in Rust Libraries

### v4

@@ III-A Sampling
[…]
@@ III-B Implementation
For our tool to function, Clang [26] must be set as the default C and C++ compiler and configured to emit LLVM bitcode. Otherwise, MiriLLI appears identical to Miri. Before executing a program, we eagerly link all LLVM bitcode files reachable from the current directory into a single module. When MiriLLI encounters a foreign function call, it looks for a corresponding function definition in the LLVM module. If we find a definition, we pass the function’s arguments through a translation layer, which lowers them into the representation used by LLI. After translation, the current Rust thread is set to join on a new LLVM thread that executes the function with the converted arguments. After the LLVM thread terminates, its return value is passed back through the translation layer and given to the Rust thread, which continues to execute. A similar process is used when an LLVM thread calls a Rust function. Though Miri is single-threaded, it supports concurrency by non-deterministically stepping through multiple simulated “threads” of execution. LLI did not originally support any form of multithreading, but we modified it to be compatible with Miri’s implementation. We support calling foreign functions from parallel Rust threads, but we do not support multithreading within LLVM.
Nearly all of Miri’s mechanisms for detecting undefined behavior are built into the core functions of the interpreter and do not require additional instrumentation of the source program. However, to detect aliasing violations, Miri inserts retag instructions wherever a reference-type value is created, passed to a function, or cast into a raw pointer. A retag produces a new permission and adds it to the stack or tree, depending on which aliasing model is in use. We did not need to insert these instructions into the LLVM IR of foreign functions, since we treat LLVM pointers as equivalent to Rust’s raw pointers. Stacked Borrows only applies a retag to raw pointers at the point where they are cast from reference-type values—which can only ever occur in Rust. Tree Borrows does not retag raw pointers at all. This allowed us to extend Miri’s error-detecting capabilities “for free” by replacing every operation that LLI uses to access memory with a foreign function call into Miri’s implementation of an equivalent operation. However, to increase test coverage, we had to address two additional difficulties.
@@
[…]
Rust’s MIR uses different calling conventions and represents values differently than LLVM IR, so we could not rely on a function’s definition to have the same signature as its binding. However, foreign function bindings are also allowed to be incorrect, so we could not wholly trust type information to guide our conversion, either. Additionally, neither Rust not LLVM have a formal model of their ABI that we could use as an oracle. When implementing our value translation layer, we settled on a conservative approach that supports the most common ABI differences while still being able to detect certain instances of incorrect bindings. We maintain the invariant that the size of a typed value on one side of the boundary must be equal to the size of the type expected by the other side. We provide a formal model of our value conversion layer in Section 2 of the supplement. It demonstrates that conversion will fail when a difference in size is detected, which is reported as undefined behavior.
[…]
Products lack a canonical form in Rust and LLVM, so we do not support passing them as variadic arguments to LLVM or as arguments from LLVM to Miri’s shims for system calls, which lack type information to guide our conversion. When an opaque pointer from LLVM is passed to one of Miri’s shims, we assign it the type *mut u8. However, if it points to stack or static memory with a size equal to that of a primitive type, we assign it to point to that type. This may lead to false positive alignment errors when we cannot resolve an underlying type and Rust expects a type other than u8. We also do not support passing aggregates by value to shims and variadic functions, since we lack a source of type information to determine the underlying layout of the struct.
@@
However, LLVM will read uninitialized memory for the purpose of propogating undef and poison values, which represent the results of indeterminate or invalid computation.
@@
We implemented two “memory modes” that allow us to identify true positive uninitialized reads of Rust-allocated memory in LLVM.
The “Uninit” mode allows load operations to read uninitialized bytes in LLVM, while the “Zeroed” mode treats all uninitialized reads as errors but zero-initializes LLVM-allocated stack and heap memory.
Tests that read uninitialized memory must be run under each mode to ensure a complete evaluation.
We cannot detect true positive uninitialized reads that occur outside of interactions with Rust, but these errors can be found with other tools [22, 29] and are out-of-scope for our evaluation.
@@ III-C Evaluation
We used MiriLLI to execute each of the viable test cases that we identified in Section III-A. We overrode each crate to use version 1.74 (nightly-2023-09-25) of the Rust toolchain with the latest compatible version of Miri (1a82975). We set the global C and C++ compilers to version 16 of Clang with optimizations disabled. We collected data using Amazon EC2 on-demand instances provisioned through CloudBank [30]. We used c6a.2xlarge instances during the initial stages of data collection, but we switched to c6a.xlarge instances when using MiriLLI. It is single-threaded, so fewer resources were necessary. All commands were executed in Docker container running Ubuntu 23.04.
@@
We executed each viable test natively and in MiriLLI under both Stacked Borrows and Tree Borrows. We configured Tree Borrows to treated all values of types derived from Unique as having the semantics of a mutable borrow. Without setting this configuration option, only Unique values in the interior of Box<T> are treated this way. We disabled isolation to allow executing non-deterministic operations and enabled symbolic alignment checking. By default, Miri will report an error if the address of a pointer is not a multiple of the requested alignment of the type of value being read or written through it. However, the base address of an allocation is not guaranteed to be a multiple of the requested alignment, so it is possible for a misaligned pointer to be “aligned” by chance. Symbolic checking avoids these false negatives by ensuring that the pointer’s offset from the base address is a multiple of the alignment of the value being read, and that the alignment of the allocation is greater than the alignment of the value. We ignored unaligned accesses in LLVM to prioritize detecting Rust-specific errors.
[…]
@@
We reason about bugs in terms of test outcomes, where an outcome includes the results of a single test under both aliasing models. We deduplicated outcomes based on exit codes, stack traces, and error logs to avoid filing redundant bug reports. Prior to deduplication, we modified error logs to remove unnecessary elements, such as references to specific memory addresses. We also included more or less detail in stack traces depending on the location of an error. For foreign errors, we used the subset of the stack trace up to the Rust boundary to avoid deduplicating errors which appeared to be identical but were caused by mistakes at different callsites. When an error occurred in Rust code, we only used the line where the error occurred, since few of these errors were caused by LLVM.
[…]
@@
Not all errors are undefined behavior, and not all instances of undefined behavior are readily exploitable. However, we still attempted to follow ethical vulnerability disclosure practices by reporting bugs privately via email before creating public reports. When we found a bug, we examined its crate’s Cargo.toml and the GitHub profile of its repository’s owner to find the contact information for its maintainers. If we were unsuccessful, we logged a public issue. We also logged issues if we had not received a response after 1 month or if the type of error did not appear to be exploitable. Our reports included a representative test case, the output from MiriLLI, and a minimal example where applicable. When a fix was trivial, we also filed a pull request. Section 1 of our supplement includes metadata from each library and links to our contributions.
@@ IV Results
We found new bugs from crates. Currently, of the bugs have been fixed, and none have been identified as security vulnerabilities. Bugs occurred slightly more frequently in LLVM than in Rust; occurred in LLVM, occurred in Rust, and were related to incorrect bindings. The majority () were found in crates with less than 100 average daily downloads in the 6 months prior to our snapshot of crates.io. However, we discovered Tree Borrows violations in separate crates which each had more than 10,000 average daily downloads.
We found Tree Borrows violations; they were the most common type of bug. Though they were caused by incorrect encapsulations or arguments in Rust, nearly all occurred in LLVM. Regarding RQ1, our results suggest that Miri’s methods for dynamically enforcing Rust’s aliasing model are useful for finding undefined behavior in Rust programs that call foreign functions.
[…]
@@
| Rust | Rust | 10 | 2 | 5 | 17 |
| Total: | 12 | 24 | 12 | 48 | |
The bugs we discovered fall into three categories. We describe Ownership errors in Section IV-A, which include all violations of Tree Borrows and accesses out-of-bounds. Allocation errors include both memory leaks and cross-language deallocation, which we describe in Section IV-C. We describe Typing errors in Section IV-B, which include incorrect foreign function bindings and values that were invalid due to incorrect alignment or initialization. Table I shows the number of bugs under each category, grouped by the location of the error and the location where a fix would need to be applied. We refer to each bug using a unique numerical ID corresponding to tables in Section 1 of our supplement.
@@ IV-A Ownership
We found Ownership bugs from crates. These include the Tree Borrows violations and accesses out-of-bounds.
round(6 + 4, 0)
@@
[…]
@@
In Rust, it is not possible to create a struct containing a self-reference using only safe code. Allowing this could lead to use-after-free errors depending on the order in which each field is dropped. It is possible, but non-trivial, to implement self-referential patterns using encapsulated unsafe operations [31]. Bug #0 demonstrates how a self-referential pattern can be implemented incorrectly. We found this bug in an encapsulation of a C library implementing a minimal file system for embedded applications. It used separate structs to represent the state of the file system and its configuration. The configuration held a mutable reference to the state, but both objects were contained in a parent struct representing the entire file system. Rust held a single mutable reference to the parent which became invalid when the state was mutated through the configuration by a foreign function.
Figure IV-A provides a minimal example of the bug and our fix. Lines that we removed are highlighted in and marked with a ‘‘-’’, while lines we added are highlighted in and marked with a ‘‘+’’. Tags within a circle indicate the state of the borrow tree prior to the fix, while tags within a square represent the state of the tree after the fix. When a tag is on a line, it indicates that the pointer or reference on that line has a permission within the subtree corresponding to the tag. The tables at the bottom display the state of the tree before and after the write access on line 14.
The Rust encapsulation for open initializes field buffer of the struct Alloc on line 9 with a mutable raw pointer cast from a mutable reference to the field cache. The struct is mutably borrowed again on line 10, creating a reference that is passed into the foreign function ffi::open. Each of these borrows, indicated by and , corresponds to an adjacent branch in the tree. Across the FFI, a write access occurs on line 14 using . This is a child access relative to and , so they transition to , but it is foreign relative to , so it becomes . This leads to an expired permission error on line 13 when is used for a read access in Rust.
We fixed this bug by wrapping parts of state in an UnsafeCell, which can be mutated through shared references. In Figure IV-A, we change the first field of Alloc to UnsafeCell<i32> and replace the mutable borrow assigned to a.buffer with a.cache.get(). This expression does not perform a retag, so both a and cache share the permission . Since a contains an UnsafeCell, the mutable borrow assigned to b on line 10 associates the tag with a special permission that can tolerate foreign writes. Consequentially, this permission is no longer disabled by the write through on line 14, so the read access on line 13 is valid.
@@
Expired permission errors typically occurred when pointers derived from mutable references were copied into the foreign heap. We observed this type of bug in separate libraries. The most notable examples were Bug #0 from bzip2 and Bug #0 from flate2, which are both popular compression libraries. In particular, flate2 is actively maintained by the Rust project and averaged more than 130,000 downloads per day during our observation period.
Figure IV-A shows a minimal example of the cyclic aliasing pattern used by each library. We use the same notation as Figure IV-A. The structs Stream and State carry pointers to each other. Both are encapsulated in Rust by a Compression object, which carries a heap-allocated instance of Stream. The Rust function new allocates memory for a Stream object and passes a mutable reference carrying the permission to ffi::init. This foreign function copies the pointer it receives to the Stream object into a newly-allocated State instance on line 20, creating a cyclic structure. Later, when the function compress is called, it creates an adjacent permission . Mutating through this permission on line 14 causes to become , making it undefined behavior to read using in the body of ffi:compress on line 23.
@@
Bug #0 also involved a mutable reference being copied into the foreign heap, but we needed a different approach to fix it. It occurred in the same file system library as Bug #0, which avoided creating heap allocations in Rust as a constraint of the embedded application context. Instead of using a Box<T> to allocate the equivalent of the Stream object in Figure IV-A, the Rust encapsulation passed around a mutable reference to a stack allocation. Similar to Bugs #0 and Bug #0, once the reference was created, it was passed to a foreign initialization function that copied the reference into a foreign heap allocation. After the initialization function returned, the mutable reference was moved into a RefCell<&mut T>. This performed a reborrow, creating a new permission as a child of the permission held by the foreign heap. Both of these permissions could be used interchangeably for read accesses, but not for write accesses. If the foreign, parent permission was mutated first, it would invalidate the child permission. If the child permission was mutated first, then both permissions would become , but the next mutation through the parent permission would cause the child permission to become , making future write accesses through the child permission undefined behavior.
To fix this error, we stored the mutable borrow as a RefCell<*mut T>, which prevented new child permissions from being created. However, the type T was a struct, and certain foreign functions required pointers to its members. Reborrowing is typically necessary to create a reference to a member of a struct, but this creates a new permission, which we needed to avoid. Instead, we used Rust’s addr_of_mut! macro, which creates a pointer to a field without borrowing. Combined, these two changes ensure that every memory access on each side of the boundary uses the same permission. The authors of Tree Borrows [10] found an equivalent bug in test cases for the Rust toolchain and applied the same fix.
@@
Rust’s PhantomData<T> is a ‘‘ghost type;’’ it behaves like an instance of T, but it can be created without providing a concrete value of that type. However, there is an exception: a value of type PhantomData<UnsafeCell<T>> does not behave like an UnsafeCell<T>; it cannot be mutated through immutable references [32]. In Bug #0, the type Opaque was declared as an alias for PhantomData<UnsafeCell<* mut T>>. It was used to represent a foreign types in Rust without having to declare their layout on both sides of the FFI boundary. Since Opaque did not contain a concrete instance of UnsafeCell, raw pointers derived from &Opaque received a read-only permission, leading to an insufficient permission error when they were written through in foreign code. We proposed fixing this error by removing the PhantomData, but other contributors are still discussing the correct fix. The relationship between PhantomData and UnsafeCell is not yet settled, so this pattern may be allowed in the future.
@@
We found access out-of-bounds errors in unique crates. Bugs #0 and #0 were found in two encapsulations for libdecnumber, which is part of the GCC toolchain. This library often used the following pattern to iterate over arrays of integers:
@@
None of these errors meaningfully involve Rust. However, the remaining errors were caused by Rust encapsulations that did not adequately enforce the preconditions of foreign function calls. Bug #0 involved the following API, which redirected its arguments to foreign functions based on the value of the size parameter:
For example, passing KeySize::K128 would call a foreign function which assumed that ex_key and key would each be 128 bytes long. However, this invariant was not enforced by the encapsulation; it was possible to cause an access out-of-bounds error by passing a KeySize larger than the length of either key. We did not propose a fix for this bug, as it could require significant changes to the API.
Bug #0 involved a foreign function that expected to receive a pointer to a string with 6 characters. It was allocated with CString::new(Vec::with_capacity(6)), exposed as a raw pointer using into_raw(), and correctly deallocated after the call using from_raw. However, the implementation of CString::new appends a null-terminator to the vector and then converts it into a value of type Box<[u8]> using into_boxed_slice(), which shortens the capacity of a vector to be equal to its length. The intended capacity was 6, but into_boxed_slice() shortened it to 1, since the string only contained a null-terminator. The encapsulation did not validate that the length of the CString instance was correct, leading to an access out-of-bounds when the C codebase read beyond the first character. We fixed this error by initializing the CString with a constant of the correct length.

### v5

@@ III-A Sampling
We evaluated our design on all compatible Rust libraries with test cases that called foreign functions. Rust libraries are referred to as ‘‘crates,’’ and they can be published at crates.io. We used a snapshot of the crates.io database taken on September 20th, 2023. It contained 125,804 unique crates, of which 96% (121,015) had at least one valid published version. We compiled each crate using version 1.74 (nightly-2023-09-25) of the Rust toolchain, matching the version used by MiriLLI. This was necessary to ensure that we could use an unmodified version of Miri as a control to identify tests that would normally fail due to lack of support for foreign function calls. It would have been infeasible to maintain separate versions of MiriLLI for multiple toolchains without compromising the consistency of our results, since Miri depends on APIs that changed significantly throughout our evaluation. There is no standard method for using a version of Miri compiled with one toolchain to execute programs compiled with another toolchain, and there is no guarantee that Miri would behave as expected under these conditions.
Of all valid crates, 67% (84,106) compiled without intervention. Of the crates that compiled, 36% (44,661) had unit tests and 9% (11,120) produced LLVM bitcode files, leaving 3% (3,785) of crates with both unit tests and LLVM bitcode. Then, we used an unmodified version of Miri to execute all of the unit tests from this subset of crates to determine which tests called foreign functions. Of the 88,637 tests that we identified, 53% (47,189) passed, 41% (36,766) failed, 4% (3,869) timed out after five minutes, and 1% (1,178) had been manually disabled. Tests can be disabled using the #[ignore] attribute or with conditional compilation directives.
Of the tests that failed in Miri, 63% (23,116) failed due to foreign function calls. We executed this subset under both Stacked and Tree Borrows using an initial build of our tool to determine which tests called foreign functions. Of all potentially viable test cases, 39% (9,130) called a foreign function we could execute. These tests originated from 25% (957) of the crates with both test cases and bitcode. Our sample only includes crates that statically link to foreign code in their default configuration; it does not include crates that default to dynamic linking or statically link in non-default configurations. However, the 957 crates that we identified were more than enough to find meaningful answers to our research questions. We used this sample to conduct our final evaluation.
We also used this opportunity to partially measure the distribution of crates that interact with foreign code. When we compiled each crate, we used an early lint pass to examine the abstract syntax tree and record the locations of all function declarations with a foreign ABI. We used data from this pass to identify 2,516 crates that declared a foreign function. This represents only 2.1% of all valid crates. However, 12,564 (10.4%) of all valid crates had a version that directly depended on at least one prior version of a crate that declared a foreign function binding. The maximum number of dependents for any crate that declared foreign functions was 6,411, with a mean of 11.0 dependents and a standard deviation of 171.8 dependents.
@@ III-B Implementation
We illustrate the data flow between each of the core components of MiriLLI in Figure 2. As indicated at the bottom of the Figure, MiriLLI requires Clang [26] to be set as the default C and C++ compiler and configured to emit LLVM bitcode during the build process for the library under test. Other than this requirement, using MiriLLI is identical to using an unmodified version of Miri.
When MiriLLI encounters a foreign function call in a Rust program, it looks for a corresponding definition in the LLVM module. After finding a definition, the function’s arguments are passed through a translation function, which lowers them into the representation used by LLI. There are two translation functions indicated by the arrows in the middle of Figure 2; one from Miri’s representation to LLI’s (left) and another from LLI’s representation to Miri’s (right). Miri passes the translated arguments to LLI, which creates a new thread to execute the function. Then, Miri sets the current current Rust thread to join on this new LLVM thread. After the LLVM thread terminates, its return value is passed back through the translation layer and given to the Rust thread, which continues to execute. A similar process is used when an LLVM thread calls a Rust function. Although Miri is single-threaded, it supports concurrency by non-deterministically stepping through multiple simulated ‘‘threads’’ of execution. LLI did not originally support any form of multithreading, but we modified it to be compatible with Miri’s implementation. We support calling foreign functions from parallel Rust threads, but we do not support multithreading within LLVM.
Nearly all of Miri’s mechanisms for detecting undefined behavior are built into the core functions of the interpreter and do not require additional instrumentation of the source program. However, to detect aliasing violations, Miri configures the Rust compiler to insert ‘‘retag’’ instructions, which are used by its implementations of Stacked and Tree Borrows. We did not need to insert these instructions into the LLVM IR of foreign functions, since we treat LLVM pointers as equivalent to Rust’s raw pointers.
We enabled LLI to detect undefined behavior by replacing several of its core operations with foreign function calls that use Miri to execute equivalent operations. As indicated by the labels next to ‘‘Miri’’ in Figure 2, all operations related to managing threads, accessing memory, and executing ‘‘shim’’ implementations of intrinsics and system calls are handled by Miri---for both interpreters. LLI manages its own local state for LLVM threads and implements the interpretation functions for each instruction, but Miri controls the process of taking a step for each thread. This architecture allowed us to extend Miri to detect errors in foreign function calls ‘‘for free’’ without having to reimplement its dynamic analyses within LLI. However, we had to address two additional difficulties to increase test coverage.
@@
Rust’s MIR uses different calling conventions and represents values differently than LLVM IR, so we could not rely on a function’s definition to have the same signature as its binding. However, foreign function bindings are also allowed to be incorrect, so we could not wholly trust type information to guide our conversion, either. Additionally, neither Rust nor LLVM has a formal model of its ABI that we could use as an oracle. When implementing our value translation functions, we settled on a conservative approach that supports the most common ABI differences while still being able to detect certain instances of incorrect bindings. We maintain the invariant that the size of a typed value on one side of the boundary must be equal to the size of the type expected by the other side. We provide a formal model of our value conversion functions in Section 2 of the Appendix. It demonstrates that conversion will fail when a difference in size is detected, which we report as undefined behavior.
Conversion is trivial for primitive types, since they have a canonical representation in Rust and LLVM. We also allow implicit casts between pointers and integers at foreign boundaries, which was necessary for 526 test cases from 70 crates. When aggregate types are passed by value, we require both the size and number of fields to match unless the aggregate is passed as an integer or by reference. However, we also frequently observed functions that expected individual aggregates to pass through multiple parameters, with one parameter for each field. This occurred only with homogeneous aggregates, where every field is identical and lacks padding. To accommodate this, we treat each field in a homogeneous aggregate as a separate argument if the entire aggregate cannot be passed as a single parameter. We apply this transformation only if the number of fixed parameters is greater than or equal to the number of remaining arguments plus the number of fields in the aggregate. This was necessary for 2% (156) of our test cases from 15% (140) of crates.
Products lack a canonical form between Rust and LLVM, so we do not support passing them as variadic arguments to LLVM or as arguments from LLVM to Miri’s shims for system calls. Each of these types of functions lacks the type information that we need to guide our conversion. When an opaque pointer from LLVM is passed to one of Miri’s shims, we assign it the type *mut u8. However, if it points to stack or static memory with a size equal to that of a primitive type, we assign it to point to that type. This may lead to false positive alignment errors when we cannot resolve an underlying type, and Rust expects a type other than u8.
@@
However, LLVM will read uninitialized memory for the purpose of propagating undef and poison values, which represent the result of an indeterminate or invalid computation.
@@
We implemented two modes that allow us to identify true positive uninitialized reads of Rust-allocated memory in LLVM. By default, we allow load instructions to read uninitialized bytes. However, MiriLLI can also be configured to treat uninitialized reads as errors while zero-initializing all LLVM-allocated stack and heap memory. Tests that read uninitialized memory must be run in each mode to ensure a complete evaluation.
We cannot detect true positive uninitialized reads that occur outside of interactions with Rust, but these errors can be found with other tools [22, 29] and are beyond the scope of our evaluation.
@@ III-C Evaluation
We used MiriLLI to execute each of the 9,130 viable test cases that we identified in Section III-A. The version of MiriLLI at this point was based on the same version of the Rust’s nightly toolchain (1.74.0) that we used in earlier stages of data collection. We set the global C and C++ compilers to version 16 of Clang and disabled optimizations. We collected data using Amazon EC2 on-demand instances provisioned through CloudBank [30]. We used c6a.2xlarge instances during the initial stages of data collection, but we switched to c6a.xlarge instances when using MiriLLI. It is single-threaded, so fewer resources were necessary. All commands were executed in a Docker container running Ubuntu 23.04.
@@
We executed each viable test natively and in MiriLLI under both Stacked Borrows and Tree Borrows. We configured Tree Borrows to treat values of the Unique<T> type as having the semantics of a mutable borrow. Without this configuration option, only the values of type Unique<T> within Box<T> are treated this way. We disabled isolation to allow executing non-deterministic operations and enabled symbolic alignment checking. By default, Miri will report an error if the address of a pointer is not a multiple of the requested alignment of the type of value being read or written through it. However, the base address of an allocation is not guaranteed to be a multiple of the requested alignment, so it is possible for a misaligned pointer to be ‘‘aligned’’ by chance. Symbolic checking avoids these false negatives by ensuring that the pointer’s offset from the base address is a multiple of the alignment of the value being read and that the alignment of the allocation is greater than the alignment of the value. We ignored unaligned accesses in LLVM to prioritize detecting Rust-specific errors.
On average, across each memory mode, 61% of tests terminated due to an unsupported operation, 19% passed, 10% timed out, 1% failed, and 9% had a potential bug. A ‘‘bug’’ includes both undefined behavior and other undesired behaviors, such as memory leaks. Of the 61% of unsupported operations, 56% were dynamically linked functions and inline assembly, which are out of scope. An additional 27% were atomic instructions and floating point types that LLI does not implement, 12% were due to types that are not supported by our value conversion layer (discussed in III-B), and 5% were shims that Miri does not fully implement.
@@
We reason about bugs in terms of test outcomes, where an outcome includes the results of a single test under both aliasing models. We deduplicated outcomes based on exit codes, stack traces, and error logs to avoid filing redundant bug reports. Prior to deduplication, we modified the error logs to remove unnecessary elements, such as references to specific memory addresses. We also included more or less detail in stack traces, depending on the location of an error. For foreign errors, we used the subset of the stack trace up to the Rust boundary to avoid deduplicating errors that appeared to be identical but were caused by mistakes at different callsites. When an error occurred in Rust, we only used the first line of the stack trace, since few of these errors were caused by LLVM.
After deduplication, we had 394 errors to investigate. Multiple factors lead to the number of bugs being significantly smaller than the number of deduplicated test outcomes. Our method of deduplication was conservative, we ignored errors from crates that had been unpublished since the start of our investigation, and we only reported Stacked Borrows violations that were also Tree Borrows violations. We also observed false positives related to alignment and accesses through addresses that we do not emulate (e.g., stdout).
@@
Not all errors are undefined behavior, and not all instances of undefined behavior are readily exploitable. However, we still attempted to follow ethical vulnerability disclosure practices by reporting bugs privately via email before creating public reports. When we found a bug, we examined its crate’s Cargo.toml and the GitHub profile of its repository’s owner to find a way to contact its maintainers privately. If we were unsuccessful, we logged a public issue. We also logged issues if we had not received a response after one month or if the type of error did not appear to be exploitable. Our reports included a representative test case, the output from MiriLLI, and a minimal example where applicable. When a fix was trivial, we also filed a pull request. Section 1 of our Appendix includes metadata for each library and links to our contributions.
@@ IV Results
We found 47 new bugs from 37 crates. As of July 29th, 2024, 22 of the bugs have been fixed, and none have been identified as security vulnerabilities. Bugs occurred slightly more frequently in LLVM than in Rust; 25 occurred in LLVM, 16 occurred in Rust, and 6 were caused by incorrect Rust bindings to foreign functions. The majority (34) were found in crates with less than 100 average daily downloads in the six months prior to our snapshot of crates.io. However, we discovered three Tree Borrows violations in separate crates that each had more than 10,000 average daily downloads. We have no reason to believe that any of these bugs would have been found by an unmodified version of Miri. We found several additional errors that occurred in Rust and did not meaningfully involve the FFI. These errors are excluded from the counts here, but appear in our raw dataset.
We found 90 tests from 37 crates where Stacked Borrows violations occurred in foreign code. However, 66% (59) of these tests passed or encountered an unsupported operation under Tree Borrows. This is mainly because 50% (45) had invalid range errors similar to the minimal example shown earlier in Figure II-A. We found several instances where slices of an array were cast into raw pointers, passed across the FFI, and then offset beyond the range that they initially borrowed. This is considered undefined behavior under Stacked Borrows, but it is valid under Tree Borrows. This provides our answer to RQ2; Tree Borrows accepts more real-world programs that call foreign functions than Stacked Borrows due to differences in handling pointer arithmetic.
@@
| Rust | Rust | 10 | 2 | 4 | 16 |
| Total: | 12 | 24 | 11 | 47 | |
Now, we answer RQ1 by identifying three categories of undefined or undesirable behavior that occur in Rust libraries that call foreign functions. We describe Ownership errors in Section IV-A, which include all violations of Tree Borrows and accesses out-of-bounds. We describe Typing errors in Section IV-B, which include incorrect foreign function bindings and uninitialized values. We describe Allocation errors in Section IV-C, which include both memory leaks and cross-language deallocation. Table I shows the number of bugs in each category, grouped by the location of the error and the location where a fix would need to be applied. We refer to each bug using a unique numerical ID corresponding to tables in Section 1 of the Appendix.
@@ IV-A Ownership
We found 24 Ownership bugs from 20 crates. These include 18 Tree Borrows violations and six accesses out-of-bounds.
@@
We found 10 insufficient permission errors caused by incorrectly casting immutable references to mutable raw pointers. In Bug #0, a chain of casts was used to convert &self into *mut T. This pointer had a read-only permission, so mutating this structure across the FFI triggered an insufficient permission error. There were six bugs with equivalent errors. The four remaining bugs involved foreign function bindings that incorrectly declared pointers as const instead of mut, which led developers to pass pointers derived from immutable references to these functions instead of mutable ones.
@@
In Rust, it is not possible to create a struct that contains a self-reference using only safe code. Allowing this could lead to use-after-free errors depending on the order in which each field is dropped. It is possible, but non-trivial, to implement self-referential patterns using safe encapsulations of unsafe operations [31]. Bug #0 demonstrates how a self-referential pattern can be implemented incorrectly. We found this bug in an encapsulation of a C library implementing a minimal file system for embedded applications. It used separate structs to represent the state of the file system and its configuration. The configuration held a mutable reference to the state, but both objects were contained in a parent struct representing the entire file system. Rust held a single mutable reference to the parent, which became invalid when the state was mutated through the configuration by a foreign function.
Figure IV-A provides a minimal example of the bug and our fix. Lines we removed are highlighted in and marked with ‘‘-’’, while lines we added are highlighted in and marked with a ‘‘+’’. Tags within a circle indicate the state of the tree prior to the fix, while tags within a square represent the state of the tree after the fix. When a tag is on a line, it indicates that the pointer or reference on that line has a permission within the subtree corresponding to the tag. The tables at the bottom display the state of the tree before and after the write access on line 18.
The Rust encapsulation for open initializes the field buffer of the struct Alloc on line 10 with a mutable raw pointer cast from a mutable reference to the field cache. The struct is mutably borrowed again on line 11, creating a reference that is passed into the foreign function ffi::open. Each of these borrows, indicated by and , corresponds to an adjacent branch in the tree. Across the FFI, a write access occurs on line 18 using . This is a child access relative to and , so they transition to . However, it is foreign access relative to , so this permission becomes . This leads to an expired permission error on line 14 when is used for a read access in Rust.
We fixed this bug by wrapping parts of the state in an UnsafeCell, which can be mutated through shared references. In Figure IV-A, we change the first field of Alloc to UnsafeCell<i32> and replace the mutable borrow assigned to a.buffer with a.cache.get(). This expression does not perform a retag, so both a and cache share permission . Since a contains an UnsafeCell, the mutable borrow assigned to b on line 11 associates the tag with a special permission that can tolerate foreign writes. Consequently, this permission is no longer disabled by the write through on line 18, so the read access on line 14 is valid.
@@
Expired permission errors occurred when pointers derived from mutable references were copied into the foreign heap. We observed this type of bug in five separate libraries. The most notable examples were Bug #0 from bzip2 and Bug #0 from flate2, which are both popular compression libraries. In particular, flate2 is actively maintained by the Rust project and averaged more than 130,000 downloads per day during our observation period.
Figure IV-A shows a minimal example of the cyclic aliasing pattern used by each library. We use the same notation as Figure IV-A. The structs Stream and State carry pointers to each other. Both are encapsulated in Rust by a Compression object, which carries a heap-allocated instance of Stream. The Rust function new allocates memory for a Stream object and passes a mutable reference carrying the permission to ffi::init. This foreign function copies the pointer it receives to the Stream object into a newly allocated State instance on line 22, creating a cyclic structure. Later, when the function compress is called, it creates an adjacent permission . Mutating through this permission on line 15 causes to become , making it undefined behavior to read using in the body of ffi:compress on line 25.
@@
Bug #0 also involved a mutable reference being copied into the foreign heap, but we needed a different approach to fix it. It occurred in the same file system library as Bug #0, which avoided creating heap allocations in Rust as a constraint of the embedded application context. Instead of using a Box<T> to allocate the equivalent of the Stream object in Figure IV-A, the Rust encapsulation passed around a mutable reference to a stack allocation. Similar to Bugs #0 and Bug #0, once the reference was created, it was passed to a foreign initialization function that copied the reference into a foreign heap allocation. After the initialization function returned, the mutable reference was moved into a RefCell<&mut T>. This performed a reborrow, creating a new permission as a child of the permission held by the foreign heap. Both of these permissions could be used interchangeably for read accesses but not for write accesses. If the foreign parent permission was mutated first, it would invalidate the child permission. If the child permission was mutated first, then both permissions would become , but the next mutation through the parent permission would cause the child permission to become , making future write accesses through the child permission undefined behavior.
To fix this error, we stored the mutable borrow as a RefCell<*mut T>, which prevented new child permissions from being created. However, the type T was a struct, and certain foreign functions required pointers to its members. Reborrowing is typically necessary to create a reference to a member of a struct, but this creates a new permission, which we needed to avoid. Instead, we used Rust’s addr_of_mut! macro, which creates a pointer to a location without borrowing it.666Rust 1.82.0 provided native syntax: &raw mut Together, these two changes ensure that all memory accesses on each side of the boundary use the same permission. The authors of Tree Borrows [10] found an equivalent bug in test cases for the Rust toolchain and applied the same fix.
@@
Rust’s PhantomData<T> is a ‘‘ghost type;’’ it behaves like an instance of T, but it can be created without providing a concrete value of that type. However, there is an exception: a value of type PhantomData<UnsafeCell<T>> does not behave like an UnsafeCell<T>; it cannot be mutated through immutable references [32]. In Bug #0, the type Opaque was declared as an alias for PhantomData<UnsafeCell<* mut T>>. It was used to represent foreign types in Rust without having to declare their layout on both sides of the FFI boundary. Since Opaque did not contain a concrete instance of UnsafeCell, raw pointers derived from &Opaque received a read-only permission, leading to an insufficient permission error when they were used for a write access in foreign code. We proposed fixing this error by removing PhantomData, but other contributors are still discussing the correct fix. The relationship between PhantomData and UnsafeCell is not yet settled, so this pattern may be allowed in the future.
@@
We found six accesses out-of-bounds in unique crates. Bugs #0 and #0 were found in two encapsulations for libdecnumber, which is part of the GCC toolchain. This library often used the following pattern to iterate over arrays of integers:
@@
None of these errors meaningfully involve Rust. However, the remaining errors were caused by Rust encapsulations that did not adequately enforce the preconditions of foreign function calls. Bug #0 involved the following API, which redirected its arguments to foreign functions based on the value of the parameter size:
For example, passing KeySize::K128 would call a foreign function that assumed that ex_key and key would each be 128 bytes long. However, this invariant was not enforced by the encapsulation; it was possible to cause an access out-of-bounds error by passing a KeySize larger than the length of either key. We did not propose a fix for this bug, as it could require significant changes to the API.
Bug #0 involved a foreign function that expected to receive a pointer to a string with six characters. It was allocated with CString::new(Vec::with_capacity(6)), exposed as a raw pointer using into_raw(), and correctly deallocated after the call using from_raw. However, the implementation of CString::new appends a null-terminator to the vector and then converts it into a value of type Box<[u8]> using into_boxed_slice(), which shortens the capacity of a vector to be equal to its length. The intended capacity was six, but into_boxed_slice() shortened it to one, appending a null-terminator. The encapsulation did not validate that the length of the CString instance was correct, leading to an access out-of-bounds when the C codebase read beyond the first character. We fixed this error by initializing the CString with a constant of the correct length.